OpenStudy (anonymous):

BEST ANSWER GETS THE MEDAL!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!1 Find the two products below. Compare and contrast, in complete sentences, the similarities and differences of the two. (x + 4)(x − 4) and (x + 4)(x + 4)

OpenStudy (mathstudent55):

Do you know FOIL?

OpenStudy (anonymous):

im not the best

OpenStudy (mathstudent55):

Let's go over using FOIL. When you multiply 2 binomials together, you need to multiply both terms of the first binomial by both terms of the second binomial. That's what FOIL does.

OpenStudy (anonymous):

@Nurali can you just give me the answer

OpenStudy (anonymous):

im in the middle of a test. please

OpenStudy (mathstudent55):

(x + 4)(x − 4) = x * x + x * (-4) + 4 * x + 4 * (-4) (x + 4)(x + 4) = x * x + 4 * x + x * 4 + 4 * 4 Now you need to simplify both exprressions above.

OpenStudy (anonymous):

@.Sam.

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

you tag mods to try to give you just the answer when your in the middle of a test?

OpenStudy (anonymous):

just this once. please

OpenStudy (amistre64):

Mods are here to make sure people are adhering to the Code of Conduct, not to make exceptions for it ...